OpenStudy (anonymous):

Solve the following system and show your work. x − 2y = 14 x + 3y = 9

sam (.sam.):

You need to isolate x for the first one, add both sides by 2y, then you got x, substitute x into 2nd equation and find y

OpenStudy (anonymous):

can you like show me with the numbers cuz im confused

OpenStudy (anonymous):

@AravindG

sam (.sam.):

\[x \cancel{− 2y + \color{red}{2y}}= 14 + \color{red}{2y} \\ \\ x=14+2y\] Now substitute x into 2nd equation and find y

OpenStudy (anonymous):

so 14 + 2y + 3y = 9

OpenStudy (ajprincess):

ya. nw solve for y.:)

OpenStudy (anonymous):

is y -1

OpenStudy (ajprincess):

u mean y=-1?

OpenStudy (anonymous):

yes

OpenStudy (ajprincess):

it is right:) Nw plu in that value in x=14+2y to find x.

OpenStudy (anonymous):

isn't x = -7

OpenStudy (ajprincess):

nope.:(